# A solution to Banach conjecture

### v3

@@ 1. Introduction
The proof of Theorem 1.1 is divided into two steps. We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding ellipsoid inside $K$. With this large ellipsoid at hand, we then establish the desired conclusion.
@@ 3. Proof of the main theorem
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [13, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
@@
Condition (1) can be reformulated as follows: for any $\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ and any $\zeta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we require that
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for every $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$, since, by choosing $\eta$ to be a basic direction, we have
$\phi_{\xi}(K\cap\xi^{\perp})=\phi_{\xi}\phi^{-1}_{\eta}(K\cap\eta^{\perp})$
for every $\xi\in\mathbb{S}^{n-1}$.
First, we introduce a function $\Gamma:\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\to\mathbb{R}_{\geq 0}$ defined by
$\Gamma(\xi)=\sup\Big\{\Big|\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|:\ \zeta,\eta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(\xi,\xi_{0})\Big\}.$
Then $\Gamma(\xi)$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Furthermore, $\Gamma(\xi)=\Gamma(\eta)$ whenever there exists a linear transformation between $K\cap\xi_{0}^{\perp}\cap\xi$ and $K\cap\xi_{0}^{\perp}\cap\eta$, as implied by Lemma 3.1.
Now fix $\xi$. For any $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$, select $\phi_{\zeta}\in\Phi(\zeta)$ such that
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=K\cap\zeta^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp}.$
We partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $\xi$ such that, for every $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$,
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}=g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $A_{2}$ consists of all $\xi$ for which there exists a vector $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ such that
$\phi_{\zeta}|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}\neq g|_{K\cap\xi^{\perp}\cap\xi_{0}^{\perp}}$
@@
We once more partition $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ into two subsets, $B_{1}$ and $B_{2}$. Define
$B_{1}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)=0\},\quad B_{2}:=\{\xi\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:\Gamma(\xi)>0\}.$
Our goal is to show that $B_{1}$ and $B_{2}$ are closed.
Note that, for $\xi\in A_{1}$ and $\zeta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$, we have
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
where
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=J(K\cap\xi_{0}^{\perp})\cap g^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}),$
@@
$J(K\cap\xi_{0}^{\perp})\cap\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=g^{-1}(g(J(K\cap\xi_{0}^{\perp}))\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp})\cap\phi_{\zeta}^{-1}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp}))=g\circ g^{-1}(J(K\cap\xi_{0}^{\perp})\cap K\cap\xi^{\perp}\cap\eta^{\perp}),$
@@
$\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))\cap\xi_{0}^{\perp}=J(K\cap\xi_{0}^{\perp})\cap\zeta^{\perp}.$
@@
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\cap\xi^{\perp}$. And thus, $\Gamma(\xi)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
We now focus on the set $B_{2}$. For any $\xi\in B_{2}$, and therefore $\xi\in A_{2}$, and for any $\zeta,\eta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\xi,\xi_{0})$ with $\zeta\neq\xi_{0}$ and $\zeta\neq\eta$, we have
$K\cap\xi^{\perp}\cap\xi_{0}^{\perp}=\phi_{\zeta}(K\cap\xi_{0}^{\perp})\cap\xi_{0}^{\perp},$
and, in addition,
$\phi^{-1}_{\zeta}(K\cap\xi^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\xi_{0}^{\perp})=\phi^{-1}_{\zeta}(K\cap\zeta^{\perp}\cap\eta^{\perp})=K\cap\xi_{0}^{\perp}\cap\phi^{-1}_{\zeta}(\eta^{\perp}).$
Hence, by exploiting the linearity of $\phi^{-1}_{\zeta}$, for any $y\in\phi^{-1}_{\zeta}(\eta^{\perp})$, we obtain $\langle\phi_{\zeta}(y),\eta\rangle=0$. Equivalently, this can be written as $y^{T}A^{T}_{\phi_{\zeta}}\eta=0$, which implies that $y\in(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$, where $A_{\phi_{\zeta}}$ denotes the matrix associated with the linear transformation $\phi_{\zeta}$. Therefore, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap(A^{T}_{\phi_{\zeta}}\eta)^{\perp}$. Furthermore, the set
$E_{\zeta,\xi}:=\{A^{T}_{\phi_{\zeta}}\eta:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\}$
forms an ellipse, and the set
$C_{\zeta,\xi}:=\left\{\frac{A^{T}_{\phi_{\zeta}}\eta}{|A^{T}_{\phi_{\zeta}}\eta|}:\eta\in\mathbb{S}^{n-1}\cap\mbox{span}(\xi,\xi_{0})\right\}$
gives a circle. And thus, there exists a linear transform between $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ and $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$ for any $\tilde{\eta}\in C_{\zeta,\xi}$. Finally, we have
[…]
for any $\tilde{\eta}\in C_{\zeta,\xi}$.
Note that for any other point $\tilde{\xi}\in B_{2}$, there is a circle $C_{\tilde{\zeta},\tilde{\xi}}$ such that
$\Gamma(\tilde{\xi})=\Gamma(\bar{\eta})$
for all $\bar{\eta}\in C_{\tilde{\zeta},\tilde{\xi}}$. Consequently, the function $\Gamma$ must be constant on $B_{2}$, since any two circles on $\mathbb{S}^{n-1}$ intersect. Therefore, $B_{2}$ is closed.
By the connectedness of $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we obtain either $B_{1}$ or $B_{2}$ as $\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$. If $B_{1}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Gamma(\xi)=0$ for any $\xi$, and hence Condition (1) holds.
If $B_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ holds, then $A_{2}=\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ implies that any (n2)(n-2)-dimensional sections of $K\cap\xi_{0}^{\perp}$ are all linearly dependent, because for every $\tilde{\eta}\in C_{\zeta,\xi}$ there is a linear transformation mapping $K\cap\xi^{\perp}\cap\xi_{0}^{\perp}$ onto $K\cap\xi_{0}^{\perp}\cap\tilde{\eta}^{\perp}$. Using the fact that all (n1)(n-1)-dimensional sections are linearly dependent, we then conclude that every (n2)(n-2)-dimensional section of $K$ is also linearly dependent. Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid. ∎
To establish that $E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$, it suffices to prove that $\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$ and $\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))$ coincide for any $k$-dimensional subspace ($1\leq k\leq n-2$). This leads to the following lemma, which is proved by induction.
Lemma 3.4.
  1. (1)
    For all $H\in G(n,k)$ with $1\leq k\leq n-2$, we have
    for every $\theta\in\mathbb{S}^{n-1}\cap H$ and $\xi,\eta\in H^{\perp}$.
  2. (2)
    The body $K$ is an ellipsoid.
Proof.
If $E$ is not an ellipsoid, then by Lemma 3.3, Condition (1) is satisfied for k=n2k=n-2. Now assume that Condition (1) holds for $k$.
For any $H\in G(n,k-1)$ and any $\xi\in H^{\perp}$, we have
$\mathrm{span}(H,\xi)\in G(n,k).$
Therefore, by the assumption of Condition (1), we obtain
for every $\theta\in\mathbb{S}^{n-1}\cap\,\mathrm{span}(H,\xi)$ and every $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$. In particular, this implies that
for all $\theta\in\mathbb{S}^{n-1}\cap H$ and all $\zeta,\eta\in\bigl(\mathrm{span}(H,\xi)\bigr)^{\perp}$, where
$\dim\bigl((\mathrm{span}(H,\xi))^{\perp}\bigr)=n-1-\dim(H)\geq 2.$
That Condition (1) is valid for k1k-1 follows from the fact that
$(\mathrm{span}(H,\xi_{1}))^{\perp}\cap(\mathrm{span}(H,\xi_{2}))^{\perp}\neq\emptyset$
for any $\xi_{1},\xi_{2}\in H^{\perp}$.
In particular, in the case $k=1$, for any $\theta\in\mathbb{S}^{n-1}$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$, we have
@@
By Lemmas 3.2 and 3.4, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$. In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.

### v4

@@ 1. Introduction
The proof of Theorem 1.1 is divided into three steps. We first select a John ellipsoid in $K\cap\xi_{0}^{\perp}$ and then use the linear invariance property of John ellipsoids to obtain a corresponding star body $E$ inside $K$. Secondly, we use the level set method (see [12, 16, 17] for details) to show $E$ an ellipsoid. With this large ellipsoid at hand, we then establish the desired conclusion.
@@ 3. Proof of the main theorem
Theorem 3.1 follows from the classical characterization by Busemann of ellipsoids (see [14, Theorem 3.1]). The convexity of the body $K$ results from the fact that if we have two points in the body, we can examine the plane they span (with the origin). Since the planar section is an ellipsoid by assumption, the entire line segment between $x$ and $y$ is contained in $K$.
@@
First, we introduce two functions $\Lambda:\mathbb{S}^{n-1}\to\mathbb{R}_{\geq 0}$ and $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
$\Lambda(\theta):=\sup_{\xi,\eta\in\mathbb{S}^{n-1}\cap\theta^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}$
and
$\Gamma(H):=\sup_{\theta\in\mathbb{S}^{n-1}\cap H\atop\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}.$
It is clear that
$\Gamma(H)=\sup_{\theta\in\mathbb{S}^{n-1}\cap H}\Lambda(\theta)$
and Condition (1) is equivalent to $\Lambda(\theta)=0$ for all $\theta\in\mathbb{S}^{n-1}$ or $\Gamma(H)=0$ for all $H\in G(n,n-2)$.
Moreover, $\Lambda,\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Furthermore, $\Gamma(H_{1})=\Gamma(H_{2})$ whenever there exists a linear transformation between $K\cap H_{1}$ and $K\cap H_{2}$, as implied by Lemma 3.1.
Now we partition G(n,n2)G(n,n-2) into two subsets, $A_{1}$ and $A_{2}$. The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. The set $A_{2}$ consists of all $H\in G(n-2)$ for which there exists a pair of vectors $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$ such that
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
@@
We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
$B_{1}:=\{H\in G(n,n-2):\Gamma(H)=0\},\quad B_{2}:=\{H\in G(n,n-2):\Gamma(H)>0\}.$
Our first goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
Note that, for $H\in A_{1}$ and $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, we have
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap H)$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. Then
$\phi_{\eta}(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}\circ g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
$\phi_{\eta}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp}))=\phi_{\xi}(\phi^{-1}_{\xi}(H)\cap J(K\cap\xi_{0}^{\perp})),$
@@
$\phi_{\eta}\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})),$
and thus,
$\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(H)\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))=H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$
@@
$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H.$
@@
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap H$. And thus, $\Gamma(H)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
We now focus on the set $B_{2}$. For any $H\in B_{2}$, and therefore $H\in A_{2}$, and for any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exists a linear transformation
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
Moreover, since $H\in B_{2}$, there is some $\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
$\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq\phi_{\xi}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. In particular, this implies $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\eta$ lying on the arc between $\xi$ and $\eta$.
Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi_{0}^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
Owing to the central symmetry of $K$, we introduce the family $\mathcal{L}(H)$ defined by
$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ is either a $2m$-circle or a single large circle passing through the antipodal point. Moreover, for any $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Lambda(\theta)$, denoted by $M$, is nonzero. By the definition of $\Gamma$, for every $H\in G(n,n-2)$ containing $\theta$, we have $\Gamma(H)=M$. Clearly, such an $H$ lies in $B_{2}$, and thus also in $A_{2}$. Hence, by the preceding argument and the fact that $\Gamma$ is invariant under linear transformations, the level set
$\{H\in G(n,n-2):\Gamma(H)=M\}$
definitely contains $\mathcal{L}(H)$ for some $H$ satisfying $\Gamma(H)=M$, where $\mathcal{L}(H)$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, where $\infty$ is the shape of a figure eight.
Observe that, by the compactness of G(n,n2)G(n,n-2), there exists a sufficiently small $\epsilon>0$ such that for every $c\in(M-\epsilon,M)$, the set
$\{H\in G(n,n-2):\Gamma(H)=c\}$
cannot contain $H\times\infty$, and therefore it must contain an (n1)(n-1)-dimensional manifold with antipodal points. Consequently, by the continuity of $\Gamma$, the level set
$\{H\in G(n,n-2):\Gamma(H)=M\}$
also contains an (n1)(n-1)-dimensional manifold with antipodal points. This implies that
$\{H\in G(n,n-2):\Gamma(H)=c\}\mbox{ and }\{H\in G(n,n-2):\Gamma(H)=M\}$
are the same set, and therefore
$\{H\in G(n,n-2):\Gamma(H)=M\}=G(n,n-2).$
We now show that if $B_{2}\neq\emptyset$, then in fact $B_{2}=G(n,n-2)$, which implies $A_{2}=G(n,n-2)$. By repeating the preceding argument for $\mathcal{L}(H)$ with any $H\in G(n,n-2)$, we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent. Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid. ∎
In particular, Condition (1) in Lemma 3.3 gives
@@
By Lemmas 3.2 and 3.3, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$. In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.